OpenStudy (anonymous):

What is the final balance for the investment? $20,000 for 3 years at 5% compounded annually

OpenStudy (anonymous):

@Whitemonsterbunny17 @quirkymalik @HelpBlahBlahBlah @helpme1.2 @happinessbreaksbones @halorazer

OpenStudy (anonymous):

@forgottensoul @wagnermar @satellite73 @coolsday @christianfilms11 @gabylovesu @HMT

OpenStudy (anonymous):

Do you know how to do the interest equation?

OpenStudy (anonymous):

no i need some one to help me like to show me so i can do it with them or some thing

OpenStudy (anonymous):

@Mindblast3r i did not get that like on how to do that:/

OpenStudy (anonymous):

\[A=P(1+\frac{ r }{ n }) \] to the n * t.

OpenStudy (mindblast3r):

So what do you want the answer, or you want to learn how to do it on your own?

OpenStudy (anonymous):

like learn how

OpenStudy (anonymous):

Where P is the starting amount, r is the interest rate, t is the number of years, and n is how many times you do it per year.

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ P=Principal(\textit{original amount, 20,000})\\ r=rate=3\%\implies \frac{3}{100}\implies 0.03\\ n=periods(\textit{amount compounded per year, annually, once a year})\\ t=years \\ \quad \\ A=P\left(1+\frac{r}{n}\right)^{nt}\implies A=20,000\left(1+\frac{0.03}{1}\right)^{1\cdot 3}\)

OpenStudy (anonymous):

oh okay

OpenStudy (mindblast3r):

These guys are very helpful. You should give them medals.

OpenStudy (anonymous):

You have 20000 dollars. Let A=20000. Interest is 5%. So after 1 year your money will be 105% of its initial value. That is A*1.05. (105% = 1.05). After another year it will be 105% of its previous value. That is A*1.05*1.05. You can say that an investment of A dollars with interest r will be A*(1+r)^n after n years. Got it?

OpenStudy (anonymous):

yeah thanks u guys

OpenStudy (mindblast3r):

You understand now?

OpenStudy (anonymous):

yes thanks

OpenStudy (mindblast3r):

:)